OpenStudy (anonymous):

Someone please help me! :) I need somone to teach me how to work this. 3. A distance-time graph indicates than an object moves 100 m in 4 s and then remains at rest for 6 s. What is the average speed of the object? 50 m/s 25 m/s 10 m/s 100 m/s

OpenStudy (anonymous):

I don't want to be given the answer I want someone to help me work it out.

OpenStudy (anonymous):

So I can find it on my own :)

OpenStudy (anonymous):

well you know s=v*t

OpenStudy (anonymous):

so v=s/t

OpenStudy (anonymous):

No I don't :(

OpenStudy (anonymous):

suppose the object starts moving at t=0..then it moves upto 4s and doesn't move for next 6s....implies total time under consideration is 10s..now use the formula speed=distance/time=100m/10s=10m/s

OpenStudy (anonymous):

whats the speed and difference?

OpenStudy (anonymous):

I'm sorry :/ I don't want to bother you but I just really need help :(

OpenStudy (anonymous):

Speed is how far something moves in a certain amount of time. For example a car driving 100 kph is going to travel 100 kilometres in 1 hour. So how far has the object in this question moved, and how long has it taken to move that far?

OpenStudy (anonymous):

it moved 100 miles in 4 seconds?

OpenStudy (anonymous):

m is for meters not miles, so 100 meters in 4 seconds the other part to this question is that it stops and waits for another 6 seconds so what is the time in total? its 10 right, so 100 meters in 10 seconds But the question wants the speed per 1 second not per 10 seconds. So how many meters did it move on average each second?

OpenStudy (anonymous):

Could you write this in a equation? Maybe I could understand it better that way.

OpenStudy (anonymous):

Its C i took that quiz already lol.